OpenStudy (anonymous):

how do you find all of the zeros of a function: ex f(x)=x^2(x+3)

OpenStudy (anonymous):

To find the zeros of a function like the one you have, you set it equal to zero.

OpenStudy (anonymous):

So, you have: x^2(x+3) = 0

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

Then you set each equal to zero.

OpenStudy (anonymous):

so you have x^2 = 0 x + 3 = 0

OpenStudy (anonymous):

Can you solve those two?

OpenStudy (anonymous):

x=3 or x=-3?

OpenStudy (anonymous):

when you isolate your x, you subtract 3 on both sides.

OpenStudy (anonymous):

x = 0 - 3 x = -3

OpenStudy (anonymous):

But, you forgot the other equation.

OpenStudy (anonymous):

x^2 = 0

OpenStudy (anonymous):

you have to square root both sides.

OpenStudy (anonymous):

so x can equal 3 and 0

OpenStudy (anonymous):

so, your final answer: x = -3 and x = 0

OpenStudy (anonymous):

Those are your zeros. Got it?

OpenStudy (anonymous):

yes i get it thank you so much

OpenStudy (anonymous):

No problem! :)